I need help with this problem: The sum of two numbers is 15 less than twice the first number. Their difference is 5 less than twice the second number. Find the numbers.
Please help! Thanks!!

Answer by jonvaliente(64)


x+y=2x-15 (1)
x-y=2y-5 (2)
Simplifying, we get:
x-y=15 (1)
-x+3y=5 (2)
Adding the 2 equations together, we get:
2y=20
y=10
Using y=10, to solve for x in (1) we get:
x-10=15
x=25
so the numbers are 25 and 10.